Assume that your risk aversion (A) is 2 and that your utility can be described by the following function:

U = E(r) - 1/2Aσ2

What is your utility for the following security x:

βx = 1.7

rm = 0.07

rf = 0.02

σ = 0.12

Please use two decimal points (or more) in your answer, and answer in decimal form (so if your calculation shows 0.13, then answer 0.13 and not 13 as in 13 %).

Also remember that you need to use a dot and not a comma to represent your decimal separator.
Answer #1

Expected return=risk free rate+beta*(market return-risk free rate)=0.02+1.7*(0.07-0.02)=0.105

Utility=0.105-0.5*2*(0.12)^2=0.0906